# Complexity Analysis of Normalizing Constant Estimation: from Jarzynski Equality to Annealed Importance Sampling and beyond

### v1

@@ Appendix A Preliminaries (Continued)
The theories of backward stochastic integral and the Girsanov theorem are adapted from Vargas et al. (2024). Here, we include relevant results and proofs to ensure a self-contained presentation.
@@
where $[X,Y]_{\cdot}$ is the cross quadratic variation process [^3] The notation used in Karatzas and Shreve (1991) is $\left\langle\cdot,\cdot\right\rangle_{\cdot}$. We use square brackets here to avoid conflict with the notation for inner product. of the local martingale parts of $X$ and $Y$.
@@
Although rarely used in practice, the backward stochastic integral is sometimes referred to as the Hänggi-Klimontovich integral in the literature. Recall that the Itô integral is defined as the limit of Riemann sums when the leftmost point of each interval is used, while the Stratonovich integral is based on the midpoint and the backward integral uses the rightmost point. The equivalence in Eq. 17 can be justified in Karatzas and Shreve (1991, Chap. 3.3).
Lemma 15 (Continuation of Lem. 1).
@@
    If we replace the SDEs in Lem. 1 with
@@
    Define the following two SDEs from 00 to $T$:
@@
    Let $\mathbb{P}^{X^{\leftarrow}}$ and $\mathbb{P}^{Y^{\leftarrow}}$ be the path measures of $X^{\leftarrow}$ and $Y^{\leftarrow}$, respectively. From Lem. 1, we know that
@@
    We present a formal proof by considering the process $\mathrm{d}Z_{t}=\sigma\mathrm{d}B_{t}$ and $Z_{0}\sim\lambda$, the Lebesgue measure. As a result, formally $Z_{t}\sim\lambda$ for all $t$, so it can also be written as $\mathrm{d}Z_{t}=\sigma\mathrm{d}B^{\leftarrow}_{t}$, $Z_{T}\sim\lambda$. The result follows by applying Lem. 1 to $X$ and $Z$ and 1. to $Y$ and $Z$.
Remark 16.
@@ Appendix B Pseudo-codes of the Algorithms
See Algorithms 1 and 2 for the detailed implementation of the AIS and RDS algorithms, respectively.
Input: The target distribution $\pi\propto\mathrm{e}^{-V}$, smoothness parameter $\beta$, total time $T$; TI annealing schedule $\lambda_{0}>...>\lambda_{K}=0$; AIS annealing schedule $\lambda(\cdot)$ with $\lambda(0)=2\beta$, AIS time points $10=\theta_{0}<...<\theta_{M}=$ 1.
@@
3// Phase 1: estimate $Z_{0}$ via TI.
@@
5Define $V_{0}:=V+\beta\|\cdot\|^{2}$, $\rho_{k}:\propto\exp\left(-V_{0}-\frac{\lambda_{k}}{2}\|\cdot\|^{2}\right)$, and $g_{k}:=\exp\left(\frac{\lambda_{k}-\lambda_{k+1}}{2}\|\cdot\|^{2}\right)$, for $k\in\left[\!\left[0,K-1\right]\!\right]$;
6
7Initialize $\widehat{Z}_{0}\leftarrow\exp\left(-V_{0}(0)+\frac{\|\nabla V_{0}(0)\|^{2}}{2(3\beta+\lambda_{0})}\right)\left(\frac{2\pi}{3\beta+\lambda_{0}}\right)^{\frac{d}{2}}$;
@@
9for $k=0$ to $K-1$ do
10       Obtain $N$ i.i.d. approximate samples $x_{1}^{(k)},...,x_{N}^{(k)}$ from $\rho_{k}$ (e.g., using LMC or proximal sampler);
@@
15
16// Phase 2: estimate $Z$ via AIS.
17
18Approximately sample $x_{0}$ from $\pi_{0}$ (e.g., using LMC or proximal sampler);
19
20Initialize $W\leftarrow-\frac{1}{2}(\lambda(\theta_{0})-\lambda(\theta_{1}))\|x_{0}\|^{2}$;
21
22for $\ell=1$ to $M-1$ do
24
26
28        // see Lem. 30 for the derivation
29
31
32 end for
33
34return $\widehat{Z}=\widehat{Z}_{0}\mathrm{e}^{-W}$
Algorithm 1 Normalizing constant estimation via AIS.
1
2Sample $X_{0}\sim\operatorname{\mathcal{N}}\left(0,I\right)$, and initialize $W:=-\frac{\|X_{0}\|^{2}}{2}-\frac{d}{2}\log 2\pi$;
3
4for $k=0$ to $N-1$ do
6
8        // see Lem. 31 for the derivation
9
11        // see Lem. 31 for the derivation
12
13 end for
14
15Update $W\leftarrow W+V(X_{t_{N}})+(T-\delta)d$;
@@ 23       Sample an independent $\xi\sim\operatorname{\mathcal{N}}\left(0,I_{d}\right)$;
17return $\widehat{Z}=\mathrm{e}^{-W}$.
@@
Appendix C Proofs for Sec. 4
C.1 A Complete Proof of Thm. 5
@@
By Girsanov theorem (Lem. 15), we have
@@
Due to Eq. 17, it suffices to calculate $\left[a(X),X\right]_{T}$. By Itô’s formula, we have
@@
Taking the integral, we immediately obtain Eq. 8, and the proof is complete. ∎
C.2 Proof of Thm. 6
@@
The proof builds on the techniques developed in Guo et al. (2025, Thm. 1). We define $\mathbb{P}$ as the path measure of the following SDE:
@@
where the vector field $(v_{t})_{t\in[0,T]}$ is chosen such that $X_{t}\sim\widetilde{\pi}_{t}$ under $\mathbb{P}$ for all $t\in[0,T]$. According to the Fokker-Planck equation [^4] We assume the existence of a unique curve of probability measures solving the Fokker-Planck equation given the drift and diffusion terms, guaranteed under mild regularity conditions (Le Bris and Lions, 2008)., $(v_{t})_{t\in[0,T]}$ must satisfy the PDE
@@
which means that $(v_{t})_{t\in[0,T]}$ generates $(\widetilde{\pi}_{t})_{t\in[0,T]}$. The Nelson’s relation (Lem. 12) implies an equivalent definition of $\mathbb{P}$ as the path measure of
@@
By Girsanov theorem (Lems. 1 and 15), it is straightforward to see that
@@
Leveraging the relation between metric derivative and continuity equation (Lem. 2), among all vector fields $(v_{t})_{t\in[0,T]}$ that generate $(\widetilde{\pi}_{t})_{t\in[0,T]}$, we can choose the one that minimizes $\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}$, thereby making $\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}=|\dot{\widetilde{\pi}}|_{t}$, the metric derivative. With the reparameterization $\widetilde{\pi}_{t}=\pi_{t/T}$, we have the following relation by chain rule:
@@
Therefore, it suffices to choose $T=\frac{32\mathcal{A}}{\varepsilon^{2}}$ to make the r.h.s. of Eq. 20 less than $\frac{1}{4}$. ∎
Appendix D Proof of Thm. 8
With the forward and backward path measures $\mathbb{P}^{\to}$ and $\mathbb{P}^{\leftarrow}$ defined in Eqs. 9 and 10, we further define the reference path measure
@@
Also, recall that the sampling path measure $\widehat{\mathbb{P}}^{\to}$ is defined in Eq. 13 starts at $\widehat{\pi}_{0}$, the distribution of an approximate sample of $\pi_{0}$. For convenience, we define the following path measure, which differs from $\widehat{\mathbb{P}}^{\to}$ only from the initial distribution:
¯$\overline{\mathbb{P}}^{\to}(x_{0:M})=\pi_{0}(x_{0})\prod_{\ell=1}^{M}\widehat{F}_{\ell}(x_{\ell-1},x_{\ell}).$ (24)
Equipped with these definitions, we first prove a lemma about a necessary condition for the estimator $\widehat{Z}$ to satisfy the desired accuracy Eq. 5.
@@
Define the estimator $\widehat{Z}:=\widehat{Z}_{0}\mathrm{e}^{-W(x_{0:M})}$, where $x_{0:M}\sim\widehat{\mathbb{P}}^{\to}$, and $\widehat{Z}_{0}$ is independent of $x_{0:M}$. To make $\widehat{Z}$ satisfy the criterion Eq. 5, it suffices to meet the following four conditions:
@@
Recall that $Z=Z_{0}\mathrm{e}^{-\Delta F}$. Using Lem. 27, we have
@@
The first term is $\leq\frac{1}{8}$ if Eq. 25 holds. To bound the second term, using the definition of TV distance and the triangle inequality, we have
^$\widehat{\mathbb{P}}^{\to}\left(\left|\frac{\mathrm{e}^{-W}}{\mathrm{e}^{-\Delta F}}-1\right|\geq\frac{\varepsilon}{8}\right)$
@@
Recall that by definition (Eqs. 13 and 24), the distributions of $x_{1:M}$ conditional on $x_{0}$ are the same under $\widehat{\mathbb{P}}^{\to}$ and $\overline{\mathbb{P}}^{\to}$. Hence, $\operatorname{TV}(\widehat{\mathbb{P}}^{\to},\overline{\mathbb{P}}^{\to})=\operatorname{TV}(\widehat{\pi}_{0},\pi_{0})$. Applying Pinsker’s inequality and leveraging Eq. 20, we have
^$\widehat{\mathbb{P}}^{\to}\left(\left|\frac{\mathrm{e}^{-W}}{\mathrm{e}^{-\Delta F}}-1\right|\geq\frac{\varepsilon}{8}\right)$
@@
Note that from Eq. 23 we know that $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\to})\leq\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})$, so if Eqs. 27, 28 and 26 hold up to some small enough absolute constants, we can achieve ^$\widehat{\mathbb{P}}^{\to}\left(\left|\frac{\mathrm{e}^{-W}}{\mathrm{e}^{-\Delta F}}-1\right|\geq\frac{\varepsilon}{8}\right)\leq\frac{1}{8}$, and therefore $\operatorname{Pr}\left(\left|\frac{\widehat{Z}}{Z}-1\right|\geq\varepsilon\right)\leq\frac{1}{4}$. ∎
In the next lemma, we show how to sample from $\pi_{0}$ and estimate $\widehat{Z}_{0}$ within the desired accuracy.
@@
  1. 1.
    Using LMC initialized at $\mu_{0}=\operatorname{\mathcal{N}}\left(0,\beta^{-1}I\right)$, the oracle complexity for obtaining a sample following a distribution $\widehat{\pi}_{0}$ that is $O(1)$-close in TV distance to $\pi_{0}$ is $\widetilde{O}(d)$.
  2. 2.
    The oracle complexity for obtaining an estimator $\widehat{Z}_{0}$ of $Z_{0}$ such that Eq. 25 holds is $\widetilde{O}\left(\frac{d^{\frac{3}{2}}}{\varepsilon^{2}}\right)$.
Remark 20.
Since $R\lesssim\frac{1}{\sqrt{\beta}}$, for both cases the dependence on $R$ is negligible.
Proof.
  1. 1.
    The bound comes from Vempala and Wibisono (2019, Theorem 2) (see also Chewi (2022, Theorem 4.2.5)). In particular, the bound there depends on $\log\operatorname{KL}(\mu_{0}\|\pi_{0})$. We show that $\operatorname{KL}(\mu_{0}\|\pi_{0})$ has a uniform upper bound over all $R\lesssim 1$. The proof is as follows.
    Note that $\pi_{0}$’s potential $V_{0}=V+\frac{2\beta}{2}\|\cdot\|^{2}$ is $\beta$-strongly-convex and $3\beta$-smooth. Let $x^{\prime}$ be its global minimizer, which satisfies $\nabla V(x^{\prime})+2\beta x^{\prime}=0$. Recall from Assump. 2 that $\nabla V(x_{*})=0$, $\|x_{*}\|\leq R$. So we have
    $2\beta\|x^{\prime}\|=\|\nabla V(x^{\prime})-\nabla V(x_{*})\|\leq\beta\|x^{\prime}-x_{*}\|\leq\beta(\|x^{\prime}\|+R)\implies\|x^{\prime}\|\leq R.$
    Therefore,
    $\operatorname{KL}(\mu_{0}\|\pi_{0})$ $=\operatorname{\mathbb{E}}_{\mu_{0}}\left[\log\mu_{0}-\log\pi_{0}\right]$
    $=\operatorname{\mathbb{E}}_{\mu_{0}}\left[-\frac{\beta}{2}\|\cdot\|^{2}+\frac{d}{2}\log\frac{\beta}{2\pi}+V_{0}+\log Z_{0}\right]$
    $=-\frac{d}{2}+\frac{d}{2}\log\frac{d}{2\pi}+\operatorname{\mathbb{E}}_{\mu_{0}}V_{0}+\log Z_{0}.$
    By strong-convexity and smoothness,
    $\operatorname{\mathbb{E}}_{\mu_{0}}V_{0}$ $\leq\operatorname{\mathbb{E}}_{\mu_{0}}\left[V_{0}(x^{\prime})+\frac{3\beta}{2}\|\cdot-x^{\prime}\|^{2}\right]=V_{0}(x^{\prime})+\frac{3\beta}{2}\left(\frac{d}{\beta}+R^{2}\right);$
    $\log Z_{0}$ $=\log\int\mathrm{e}^{-V_{0}(x)}\mathrm{d}x\leq\log\int\exp\left(-V_{0}(x^{\prime})-\frac{\beta}{2}\|x-x^{\prime}\|^{2}\right)\mathrm{d}x$
    $=-V_{0}(x^{\prime})+\frac{d}{2}\log\frac{\beta}{2\pi},$
    so we conclude that
    $\operatorname{KL}(\mu_{0}\|\pi_{0})\leq d+d\log\frac{\beta}{2\pi}+\frac{3\beta R^{2}}{2}.$
  2. 2.
    The result is adapted from Ge et al. (2020, Section 3), with two key modifications. First, we relax their assumption that the global minimizer is at zero, requiring instead that the global minimizer $x^{\prime}$ satisfies $\|x^{\prime}\|\leq R\lesssim\frac{1}{\sqrt{\beta}}$. Second, we use replace their Metropolis-Hasting adjusted Langevin algorithm (MALA) with the proximal sampler (Fan et al., 2023), which achieves improved dimensional dependence. For completeness, we include a proof sketch in Sec. H.2 and refer the readers to the original work for full technical details. Our analysis confirms that these relaxations have negligible impact on the final bounds.
Next, we study how to satisfy the conditions in Eqs. 27 and 28 while minimizing oracle complexity. Given that we already have an approximate sample from $\pi_{0}$ and an accurate estimate of $Z_{0}$, we proceed to the next step of the AIS algorithm. Since each transition kernel requires one call to the oracle of $\nabla V$, and by plugging in $f_{\theta}\leftarrow V+\frac{\lambda(\theta)}{2}\|\cdot\|^{2}$ in AIS (Thm. 7), the work function $W(x_{0:M})$ is independent of $V$, it follows that the remaining oracle complexity is $M$. The result is formalized in the following lemma.
Lemma 21.
To minimize the oracle complexity, it suffices to find the minimal $M$ such that there exists a sequence $10=\theta_{0}<\theta_{1}<...<\theta_{M}=$ 1 satisfying the following three constraints:
@@
where the infimum is taken among all path measures from 00 to $T_{\ell}$ with the marginal constraints $\mathbf{P}^{\ell}_{0}=\pi_{\theta_{\ell-1}}$ and $\mathbf{P}^{\ell}_{T_{\ell}}=\pi_{\theta_{\ell}}$; $\mathbf{Q}^{\ell}$ is the path measure of Eq. 12 (i.e., LD with target distribution $\pi_{\theta_{\ell}}$) initialized at $X_{0}\sim\pi_{\theta_{\ell}}$.

### v2

@@ A.1 Stochastic Analysis: Forward-backward SDEs and Girsanov Theorem
We now present the theory of backward stochastic integral and the Girsanov theorem, which are adapted from Vargas et al. (2024). Here, we include relevant results and proofs to ensure a self-contained presentation.
The backward SDE can be perceived as the time-reversal of a forward SDE:
Definition 1 (Backward SDE).
Given a BM $(B_{t})_{t\in[0,T]}$, let its time-reversal be $(B^{\leftarrow}_{t}:=B_{T-t})_{t\in[0,T]}$. We say that a process $(X^{\leftarrow}_{t})_{t\in[0,T]}$ satisfies the backward SDE
The forward and backward SDEs are related through the following Nelson’s relation:
@@
We now define the concept of backward stochastic integral, which allows us to represent the RN derivative between path measures of forward and backward SDEs.
@@
where $X,Y]_{\cdot}$ is the cross quadratic variation process [^5] The notation used in Karatzas and Shreve (1991) is $\left\langle\cdot,\cdot\right\rangle_{\cdot}$. We use square brackets here to avoid conflict with the notation for inner product. of the local martingale parts of $X$ and $Y$.
@@
Although rarely used in practice, the backward stochastic integral is sometimes referred to as the Hänggi-Klimontovich integral in the literature. Recall that the Itô integral is defined as the limit of Riemann sums when the leftmost point of each interval is used, while the Stratonovich integral is based on the midpoint and the backward integral uses the rightmost point. The equivalence in ˜15 can be justified in Karatzas and Shreve (1991, Chap. 3.3).
Lemma 4 (Continuation of Lem.˜2).
@@
    If we replace the SDEs in Lem.˜2 with
@@
    Define the following two SDEs from 0 to $T$:
@@
    Let $\mathbb{P}^{X^{\leftarrow}}$ and $\mathbb{P}^{Y^{\leftarrow}}$ be the path measures of $X^{\leftarrow}$ and $Y^{\leftarrow}$, respectively. From Lem.˜2, we know that
@@
    We present a formal proof by considering the process $\mathrm{d}Z_{t}=\sigma\mathrm{d}B_{t}$ and $Z_{0}\sim\lambda$, the Lebesgue measure. As a result, formally $Z_{t}\sim\lambda$ for all $t$, so it can also be written as $\mathrm{d}Z_{t}=\sigma\mathrm{d}B^{\leftarrow}_{t}$, $Z_{T}\sim\lambda$. The result follows by applying Lem.˜2 to $X$ and $Z$ and 1. to $Y$ and $Z$.
A.2 Optimal Transport: Wasserstein Geometry and Metric Derivative
For two probability measures $\mu,\nu$ on $\mathbb{R}^{d}$ with finite second-order moments, the Wasserstein-2 ($\textbf{W}_{\textbf{2}}$) distance between $\mu$ and $\nu$ is defined as $W_{2}(\mu,\nu)=\inf_{\gamma\in\Pi(\mu,\nu)}\left(\int\|x-y\|^{2}\gamma(\mathrm{d}x,\mathrm{d}y)\right)^{1/2}$, where $\Pi(\mu,\nu)$ is the set of all couplings of $(\mu,\nu)$. The Brenier’s theorem states that when $\mu$ has a Lebesgue density, then there exists a unique coupling $\left(\operatorname{id}\times T_{\mu\to\nu}\right)_{\sharp}\mu$ that reaches the infimum. Here, $\sharp$ stands for the push-forward of a measure (defined by $T_{\sharp}\mu(\cdot)=\mu(\{\omega:T(\omega)\in\cdot\})$), and $T_{\mu\to\nu}$ is known as the OT map from $\mu$ to $\nu$, which can be written as the gradient of a convex function.
Lemma 5.
Finally, we define the action of an AC curve of probability measures $(\rho_{t})_{t\in[a,b]}$ as $\int_{a}^{b}|\dot{\rho}|_{t}^{2}\mathrm{d}t$, which plays a key role in characterizing the efficiency of a curve for normalizing constant estimation. For basic properties of the action and its relation to isoperimetric inequalities such as log-Sobolev and Poincaré inequalities (see definitions below), we refer the reader to Guo et al. (2025, Lem. 3 & Ex. 1).
@@ Appendix B Pseudo-codes of the Algorithms
See Algorithms˜1 and 2 for the detailed implementation of the AIS and RDS algorithms, respectively.
@@
Input: The target distribution $\pi\propto\mathrm{e}^{-V}$, smoothness parameter $\beta$, total time $T$; TI number of intermediate distributions $K$, annealing schedule $\lambda_{0}>...>\lambda_{K}=0$, number of particles NN; AIS steps $M$, annealing schedule $\lambda(\cdot)$ with $\lambda(0)=2\beta$, time points $0=\theta_{0}<...<\theta_{M}=1$.
2
3
4// Phase 1: estimate $Z_{0}$ via TI.
5
6Define $V_{0}:=V+\beta\|\cdot\|^{2}$, $\rho_{k}:\propto\exp\left(-V_{0}-\frac{\lambda_{k}}{2}\|\cdot\|^{2}\right)$, and $g_{k}:=\exp\left(\frac{\lambda_{k}-\lambda_{k+1}}{2}\|\cdot\|^{2}\right)$, for $k\in\left[\!\left[0,K-1\right]\!\right]$;
7
8Initialize $\widehat{Z}_{0}\leftarrow\exp\left(-V_{0}(0)+\frac{\|\nabla V_{0}(0)\|^{2}}{2(3\beta+\lambda_{0})}\right)\left(\frac{2\pi}{3\beta+\lambda_{0}}\right)^{\frac{d}{2}}$;
9
10for $k=0$ to K1K-1 do
11   Obtain NN i.i.d. approximate samples $x_{1}^{(k)},...,x_{N}^{(k)}$ from $\rho_{k}$ (e.g., using LMC or proximal sampler);
12
14
15 end for
16
17// Phase 2: estimate $Z$ via AIS.
18
19Approximately sample $x_{0}$ from $\pi_{0}$ (e.g., using LMC or proximal sampler);
20
21Initialize $W\leftarrow-\frac{1}{2}(\lambda(\theta_{0})-\lambda(\theta_{1}))\|x_{0}\|^{2}$;
22
23for $\ell=1$ to M1M-1 do
25
27
29 // see Lem.˜12 for the derivation.
30
32
33 end for
34
35return $\widehat{Z}=\widehat{Z}_{0}\mathrm{e}^{-W}$
Algorithm 1 Normalizing constant estimation via AIS.
1
@@ 24   Sample an independent $\xi\sim\operatorname{\mathcal{N}}\left(0,I_{d}\right)$;
3Sample $X_{0}\sim\operatorname{\mathcal{N}}\left(0,I\right)$, and initialize $W:=-\frac{\|X_{0}\|^{2}}{2}-\frac{d}{2}\log 2\pi$;
@@
5for $k=0$ to N1N-1 do
7 // this can be done by sampling $\xi_{1},\widetilde{\xi}_{2}\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\operatorname{\mathcal{N}}\left(0,I\right)$ and setting $\xi_{2}=\rho_{k}\xi_{1}+\sqrt{1-\rho_{k}^{2}}\widetilde{\xi_{2}}$
@@
10 // see Lem.˜13 for the derivation
@@
17Update $W\leftarrow W+V(X_{t_{N}})-(T-\delta)d$;
18
19return $\widehat{Z}=\mathrm{e}^{-W}$.
@@
Appendix C Proofs for Sec.˜3
C.1 A Complete Proof of Thm.˜1
@@
By Girsanov theorem (Lem.˜4), we have
@@
Due to ˜15, it suffices to calculate $\left[a(X),X\right]_{T}$. By Itô’s formula, we have
@@
Taking the integral, we immediately obtain ˜4, and the proof is complete. ∎
C.2 Proof of Thm.˜2
@@
The proof builds on the techniques developed in Guo et al. (2025, Thm. 1), yet with new components including backward SDE and the corresponding version of the Girsanov theorem.
We define $\mathbb{P}$ as the path measure of the following SDE:
@@
where the vector field $(v_{t})_{t\in[0,T]}$ is chosen such that $X_{t}\sim\widetilde{\pi}_{t}$ under $\mathbb{P}$ for all $t\in[0,T]$. According to the Fokker-Planck equation, [^6] We assume the existence of a unique curve of probability measures solving the Fokker-Planck equation given the drift and diffusion terms, guaranteed under mild regularity conditions (Le Bris and Lions, 2008). $(v_{t})_{t\in[0,T]}$ must satisfy the PDE
@@
which means that $(v_{t})_{t\in[0,T]}$ generates $(\widetilde{\pi}_{t})_{t\in[0,T]}$. The Nelson’s relation (Lem.˜3) implies an equivalent definition of $\mathbb{P}$ as the path measure of
@@
By Girsanov theorem (Lems.˜2 and 4), it is straightforward to see that
@@
Leveraging the relation between metric derivative and continuity equation (Lem.˜5), among all vector fields $(v_{t})_{t\in[0,T]}$ that generate $(\widetilde{\pi}_{t})_{t\in[0,T]}$, we can choose the one that minimizes $\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}$, thereby making $\|v_{t}\|_{L^{2}(\widetilde{\pi}_{t})}=|\dot{\widetilde{\pi}}|_{t}$, the metric derivative. With the reparameterization $\widetilde{\pi}_{t}=\pi_{t/T}$, we have the following relation by chain rule:
@@
Therefore, it suffices to choose $T=\frac{32\mathcal{A}}{\varepsilon^{2}}$ to make the r.h.s. of ˜18 less than $\frac{1}{4}$. ∎
Appendix D Proof of Thm.˜4
With the forward and backward path measures $\mathbb{P}^{\to}$ and $\mathbb{P}^{\leftarrow}$ defined in ˜5 and 6, we further define the reference path measure
@@
Also, recall that the sampling path measure $\widehat{\mathbb{P}}^{\to}$ is defined in ˜9 starts at $\widehat{\pi}_{0}$, the distribution of an approximate sample of $\pi_{0}$. For convenience, we define the following path measure, which differs from $\widehat{\mathbb{P}}^{\to}$ only from the initial distribution:
$\overline{\mathbb{P}}^{\to}(x_{0:M})=\pi_{0}(x_{0})\prod_{\ell=1}^{M}\widehat{F}_{\ell}(x_{\ell-1},x_{\ell}).$ (22)
Equipped with these definitions, we first prove a lemma about a necessary condition for the estimator $\widehat{Z}$ to satisfy the desired accuracy ˜1.
@@
Define the estimator $\widehat{Z}:=\widehat{Z}_{0}\mathrm{e}^{-W(x_{0:M})}$, where $x_{0:M}\sim\widehat{\mathbb{P}}^{\to}$, and $\widehat{Z}_{0}$ is independent of $x_{0:M}$. To make $\widehat{Z}$ satisfy the criterion ˜1, it suffices to meet the following four conditions:
@@
Recall that $Z=Z_{0}\mathrm{e}^{-\Delta F}$. Using Lem.˜9, we have
@@
The first term is $\leq\frac{1}{8}$ if ˜23 holds. To bound the second term, using the definition of TV distance and the triangle inequality, we have
$\widehat{\mathbb{P}}^{\to}\left(\left|\frac{\mathrm{e}^{-W}}{\mathrm{e}^{-\Delta F}}-1\right|\geq\frac{\varepsilon}{8}\right)$
@@
Recall that by definition ˜9 and 22, the distributions of $x_{1:M}$ conditional on $x_{0}$ are the same under $\widehat{\mathbb{P}}^{\to}$ and $\overline{\mathbb{P}}^{\to}$. Hence, $\operatorname{TV}(\widehat{\mathbb{P}}^{\to},\overline{\mathbb{P}}^{\to})=\operatorname{TV}(\widehat{\pi}_{0},\pi_{0})$. Applying Pinsker’s inequality and leveraging ˜18, we have
$\widehat{\mathbb{P}}^{\to}\left(\left|\frac{\mathrm{e}^{-W}}{\mathrm{e}^{-\Delta F}}-1\right|\geq\frac{\varepsilon}{8}\right)$
@@
Note that from ˜21 we know that $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\to})\leq\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})$, so if ˜25, 26 and 24 hold up to some small enough absolute constants, we can achieve $\widehat{\mathbb{P}}^{\to}\left(\left|\frac{\mathrm{e}^{-W}}{\mathrm{e}^{-\Delta F}}-1\right|\geq\frac{\varepsilon}{8}\right)\leq\frac{1}{8}$, and therefore $\operatorname{Pr}\left(\left|\frac{\widehat{Z}}{Z}-1\right|\geq\varepsilon\right)\leq\frac{1}{4}$. ∎
Remark 5.
As the TV distance is always upper bounded by 11, one can in fact write ˜24, 25 and 26 in a more precise way: $\operatorname{TV}(\widehat{\pi}_{0},\pi_{0})\leq 2^{-5}$, $\operatorname{KL}(\mathbb{P}\|\mathbb{P}^{\leftarrow})\leq 2^{-13}\varepsilon^{2}$, and $\operatorname{KL}(\mathbb{P}\|\overline{\mathbb{P}}^{\to})\leq 2^{-8}$.
Next, we study how to satisfy the conditions in ˜25 and 26 while minimizing oracle complexity. Given that we already have an approximate sample from $\pi_{0}$ and an accurate estimate of $Z_{0}$, we proceed to the next step of the AIS algorithm. Since each transition kernel requires one call to the oracle of $\nabla V$, and by plugging in $f_{\theta}\leftarrow V+\frac{\lambda(\theta)}{2}\|\cdot\|^{2}$ in AIS (Thm.˜3), the work function $W(x_{0:M})$ is independent of $V$, it follows that the remaining oracle complexity is $M$. The result is formalized in the following lemma.
@@
To minimize the oracle complexity, it suffices to find the minimal $M$ such that there exists a sequence $0=\theta_{0}<\theta_{1}<...<\theta_{M}=1$ satisfying the following three constraints:
@@
where the infimum is taken among all path measures from 0 to $T_{\ell}$ with the marginal constraints $\mathbf{P}^{\ell}_{0}=\pi_{\theta_{\ell-1}}$ and $\mathbf{P}^{\ell}_{T_{\ell}}=\pi_{\theta_{\ell}}$; $\mathbf{Q}^{\ell}$ is the path measure of ˜8 (i.e., LD with target distribution $\pi_{\theta_{\ell}}$) initialized at $X_{0}\sim\pi_{\theta_{\ell}}$.